### v1

# Call Me When Necessary: LLMs can Efficiently and Faithfully Reason over Structured Environments

Abstract
Large Language Models (LLMs) have shown potential in reasoning over structured environments, e.g., knowledge graphs and tables. Such tasks typically require multi-hop reasoning, i.e., match natural language utterance with instances in the environment. Previous works adopt LLMs to incrementally build a reasoning path, where LLMs either invoke tools or pick up items by step-by-step interacting with the environment. We propose Reasoning-Path-Editing (Readi), a novel framework where LLMs can efficiently and faithfully reason over structured environments. In Readi, LLMs initially generate a reasoning path given a query, and edit the path only when necessary. We instantiate the path on structured environments and provide feedback to edit the path if anything goes wrong. Experimental results on three KGQA and two TableQA datasets show the effectiveness of Readi, significantly surpassing all LLM-based methods (by Hit@1 9.1% on WebQSP, 12.4% on MQA-3H and 9.5% on WTQ), comparable with state-of-the-art fine-tuned methods (67% on CWQ and 74.7% on WebQSP) and substantially boosting the vanilla LLMs (by 14.9% on CWQ). Our code will be available upon publication.
Call Me When Necessary: LLMs can Efficiently and Faithfully Reason over Structured Environments
[…]
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance in NLP fields Ouyang et al. (2022); OpenAI (2023); Liang et al. (2023). To further unleash their reasoning ability in complex scenarios, delicate strategies are proposed to equip LLMs with human-like thought process (e.g., Chain-of-thought Wei et al., 2022) or leverage them as autonomous agents capable of planning, reflection and execution Yao et al. (2023); Liu et al. (2024b). One compelling scenario where LLMs showcase their potential is reasoning over structured environments (SEs) Jiang et al. (2023a); Sun et al. (2024). With dedicate schemas, SEs (e.g., knowledge graphs, tables) abstract real-world semantics for representing, storing, and querying data with relational structures. For instance, Freebase Bollacker et al. (2008) captures 45M entities and 3B facts over 100 domains, organized in triple patterns. The crux of successful reasoning lies in bridging the gap between natural language and the mechanism of how the SEs are represented and operated Gu et al. (2023); Li et al. (2023).
Figure 1: An illustration of our proposed framework, Readi, where LLMs initially generate a reasoning path, and when necessary, edit this path. We instantiate the path on structured environments and invoke editing if the instantiation gets stuck.
While LLMs exhibit promising capabilities, their performance often falls short when faced with multi-hop reasoning involving large-scale SEs. To faithfully reason, prior works adopt an iterative way that start from certain elements (e.g., entity, relation in KG, column in Table), instantiate on SEs and then gradually expand the reasoning path while consuming just relevant portions instead of the entire environment Sun et al. (2024); Jiang et al. (2023a). This step-by-step interaction could mitigate hallucination of LLMs to some extent. However, the reasoning efficiency is sacrificed and thus hinder the practical feasibility. For a simple constraint, say “the daughter of Obama”, such methods probably require two LLM-calls to first query the relations around “Obama” and then select “father_of” from returned candidates. Moreover, at each step, the LLMs makes one choice based on history, making it prone to error propagation. Alternatively, fine-tuned methods inject environments into model parameters by tuning with human-labeled supervision. During inference, they recall schema patterns to build reasoning paths Zhang et al. (2022); Saxena et al. (2020); Luo et al. (2024); Ding et al. (2024) without interaction with SEs. This end-to-end paradigm is efficient, however, it is never ensured that the model output can be grounded on SEs. Study shows that 50% paths of RoG Luo et al. (2024) failed to yield faithful results. In addition, they heavily rely on annotations, which are difficult to obtain for large-scale SEs. Further research is thus needed to achieve efficient (i.e., less LLM-calls) and faithful (i.e., grounded on SEs) reasoning over massive SEs.
Therefore, we seek to propose an interaction paradigm that leverage LLMs to support complex reasoning on large-scale SEs faithfully and efficiently. Conducting empirical study on KGQA task, we find that 46%-60% of reasoning paths initially generated by LLM can be well instantiated, which inspires us to fully exploit LLM’s intrinsic planning ability in complex reasoning. While the idea of plan-and-refine is straightforward and applied in various real world tasks Qiao et al. (2023); Collaboration (2023); Ahn et al. (2022), it’s worth noting that there exists few research on application in SEs when the initial plan encounters obstacles.
In this paper, we propose Reasoning-Path-Editing (Readi), a novel framework that leverages the intrinsic planning ability of LLMs (Figure 1). In Readi, LLMs initially generate a reasoning path, which is then instantiated on SEs to facilitate faithful reasoning. Path editing is triggered only if corrections are necessary. This way, we alleviate the burden of step-by-step interaction for LLMs, resulting in improved overall efficiency. To harness the information of large-scale SEs, instead of injecting the entire static SEs into the model, we collect the reasoning log as immediate feedback which includes details such as position of stuck points, associated relations, half-way done instances, etc. This dynamic guidance refines the reasoning path more targeted and further enhances faithfulness. Our experiments on Question Answering over Knowledge Graph (KGQA) and table (TableQA) demonstrate that Readi surpasses existing solutions in terms of both efficiency and accuracy.
We summerize our contributions as follows:
$\bullet$ We introduce Readi, a novel framework where LLMs reason efficiently and faithfully over large-scale structured environments. Notably, Readi is the first to fully harness LLMs intrinsic planning ability for reasoning in such contexts.
$\bullet$ In comprehensive experiments across five multi-hop reasoning tasks in KGQA and TableQA, Readi outperforms other LLM-based solutions and surpasses most fine-tuned methods. Specifically, it achieves 67.0% Hit@1 on CWQ, 78.7% on WebQSP and state-of-the-art results on MQA-1H.
$\bullet$ We also give detailed analysis which highlights the performance of Readi’s reasoning path generation and editing modules. Experiments demonstrate that Readi, with an average of 1.55 calls for editing, significantly reduces the number of LLM calls compared to the step-by-step interaction paradigm (which costs 4 to 8 calls). Furthermore, the reasoning log reveals that Readi exhibits characteristics akin to human thought process.
2 Related Works
Step-by-step reasoning over structured environments by LLMs. Introducing massive SEs (e.g., Freebase Bollacker et al. (2008) captures 45M entities and 3B facts over 100 domains) into LLMs context windows is not practical. Existing works break the task down to incrementally construct a reasoning path. They either treat LLMs as an agent to invoke some tools based on history states and observations  Liu et al. (2024b); Qin et al. (2023), or design specific procedures where LLMs are responsible for picking up elements on SEs Gu et al. (2023); Jiang et al. (2023a); Sun et al. (2024). These works reach faithfulness by leveraging the language processing ability of LLMs for tool or element selection. However, the performance is actually concerned with three shortcomings: 1) The iterative interaction with SEs is cumbersome, requiring quite a few LLM-calls, which is especially not efficient for complex reasoning task. 2) The greedy step-by-step decision lacks a global view of the path, making it prone to error propagation. 3) The prompts are lengthy where LLMs may lose attention of history or candidates Liu et al. (2024a). To ease these problems, we propose to directly generate a reasoning path, and edit it only when necessary, i.e., the instantiation gets stuck.
End-to-End reasoning over structured environments by fine-tuning. Fine-tuned models memorize the environments through training over annotations. They either directly generate a path and then ground it on SEs Luo et al. (2024); Huang et al. (2023b); Shu et al. (2022); Hu et al. (2022), or retrieve relevant items to build a path Zhang et al. (2022); Saxena et al. (2020); Ding et al. (2024). These end-to-end methods show efficiency with no interaction with SEs. However, they have three limitations: 1) The grounding of reasoning path only depends on model outputs, without ensuring faithfulness on SEs. To remedy this issue, they rely on a wider beam at the expense of more retrieved instances. 2) They rely heavily on annotations, which are expensive for massive environments. 3) The performance drops substantially on data unseen during training Gu et al. (2021); Huang et al. (2023a), which is common in real-world scenarios. To alleviate these problems, instead of fine-tuning, we propose to instantiate LLMs reasoning path. If anything goes wrong, we call LLMs to edit the path.
Plan-and-Refine Reasoning with LLMs. For faithfulness of LLMs reasoning, previous works adopt LLMs to refine the output Pan et al. (2023). Some require LLMs to self-correct, prompting them to identify and correct errors Madaan et al. (2023); Pourreza and Rafiei (2023). Such methods achieve limited improvement, since they rely only on the intrinsic knowledge of LLMs, without any access to the environment. Alternatively, other works require LLMs to refine the previous plan based on environmental feedback (e.g. code error messages) Chen et al. (2024); Qiao et al. (2023). The feedback provides the execution results and possible errors, which are more purposeful and thus effective. However, how to collect feedback for large-scale structured environments remains an open question. In Readi, we collect immediate reasoning log through the instantiation of reasoning path, including the position of error, half-way done instances and associated relations.
3 Task Definition
Reasoning over structured environments (SEs), e.g., question answering over knowledge graphs (KGQA) and tables (TableQA), typically requires matching a natural language question with path instances in SEs to constrain the answer. An intermediate reasoning path is a bridge between the question description and SEs. Figure 2 exemplifies some reasoning paths and their instances on KG. Note that a KG is a set of triple patterns, i.e., $\{(e,r,e^{\prime})|e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\}$, where $\mathcal{E}$ and $\mathcal{R}$ refers to the set of entity and relation, respectively.
Figure 2: Examples of the question, reasoning path, and corresponding path instances on knowledge graph.
[…]
Figure 3: A running example of Readi on KGQA. An LLM initially generate an reasoning path for a question. Then, we instantiate it on KG. If anything goes wrong (the path from “France”), we collect some error messages and call an LLM to edit the path. Finally, an LLM answers the question based on the KG instances.
Following Sun et al. (2018), we model reasoning over structured environments as an retrieve-then-reason task, where we leverage LLMs to build the reasoning path given $Q$. Then we reason over the path instances the obtain the answer.
4 Methodology
4.1 Overview
For better illustration, we adopt KGQA, a challenging scenario of multi-hop reasoning over massive environments, to showcase our concrete implementation. A running example is shown in Figure 3. Given a question and topic entities, we leverage the intrinsic planning ability of LLMs to generate initial reasoning path $P$ (Section 4.2). Then we instantiate $P$ on KG (Section 4.3). If the entire $P$ is successfully instantiated, we take the intersection of all instances then generate the answer. If any path in $P$ gets stuck, we collect error messages to guide further refinement (Section 4.4). Such process runs until $P$ is fully instantiated or a maximum edit time is reached.
4.2 Reasoning Path Generation
Inspired by Li et al. (2023), we leverage in-context learning (ICL) to generate the initial reasoning path, as shown in Figure 3(a). Given a question and $n$ topic entities, we utilize Chain-of-Thought (CoT) to generate the initial path, consisting of $n$ constraints starting from each topic entity. For the example in Figure 3, we have a two-constrained path (from “France” to constrain the bordering countries and from “Nijmegen” to airports serving it and then to countries containing these airports). Please refer to Appendix F for concrete prompts.
4.3 Reasoning Path Instantiation
We instantiate the reasoning path $P$ on KG and merge the instances. The main difficulty lies in how to sequentially match the natural language (NL) relations in $P$ with relations in KG. Our instantiation involves two steps: relation-binding and path-connecting, as shown in Figure 3(b).
For relation-binding, given a path (constraint) consisting of some NL relations (e.g., for “[France] border”, we have one relation “border” expressed in NL), we bind them with relation schemas in KG. Since a relation $\boldsymbol{r}$ in NL may have similar relation schemas in KG with analogical format and semantic meanings Li et al. (2023), we leverage BM25 and Contriever Izacard et al. (2022) to retrieve similar relations $\boldsymbol{\hat{r}}$ as candidates for $\boldsymbol{r}$. For example, we bind “border” with 3 candidate relations “adjoin, near_by, locations” in KG (the orange italicized relations in Figure 3(b). This way, we obtain candidate relations for all relations in $P$.
For path-connecting, note that the starting entity of each path is given. For the example in Figure 3(b), we need to check if there exists any path instance in KG from “France” where relations in this instance sequentially match the bound candidate relations of all relations in $P$. Specifically, if any relations in candidates “adjoin, near_by, locations” connects to “France” (the orange bold “adjoin”), we obtain entities satisfying the constraint “[France] border”, so the NL relation “border” is instantiated to “adjoin” on KG. Then, we repeat the same process for the remaining relations in $P$ to finally instantiate the whole path.
If any path in $P$ is not successfully instantiated, e.g., none of the candidate relations $\boldsymbol{\hat{r}}$ connects to the current entity, this is the “necessary” time to invoke editing (Section 4.4). If the entire $P$ is successfully instantiated or the maximum editing time is reached, we merge all KG instances by intersection to answer the question (Section 4.5). Please refer to Appendix B.1 for concrete implementation.
4.4 Reasoning Path Editing
The goal of editing is to help LLMs identify the error of previous reasoning path and provide some error messages, consisting of 2 steps: summarization and preparation, similar to an error traceback of coding, as shown in Figure 3(c).
For summarization, we categorize the reasons of error as follows: i. “irrelevant NL relation $\boldsymbol{r_{err}}$”: none of the KG relation candidates of $\boldsymbol{r_{err}}$ connects to current entity; ii. “empty reasoning path”; iii. “path ends with compound nodes” [^1] Compound value node, i.e., blank node, is typical in KG to express some complex entities, such as an event.. Thus, if the instantiation goes wrong, we detect error position $\boldsymbol{err}$, current NL relation $\boldsymbol{r_{err}}$ (none for reason iii) and current ending entity $\boldsymbol{\hat{e}_{err}}$ (before which the NL relations are successfully instantiated).
For preparation, we collect some useful messages: 1) we show the reason of error. 2) we maintain the currently halfway-done instances. 3) relations around $\boldsymbol{\hat{e}_{err}}$ is accessible, which might be the candidates to amend the halfway-done instances. Finally, an LLM is called to edit the previous path based on $Q$ and these error messages.
[…]
4.5 QA Reasoning
Upon obtaining the merged KG instances $\mathcal{S}_{q}$, i.e., the intersection of instances for each path (constraint), we build an LLM-based reasoning module to answer the question $Q$. We concatenate $Q$ and $\mathcal{S}_{q}$ as input and prompt an LLM to generate the answer by ICL. Note that the form of $\mathcal{S}_{q}$ is a set of triple patterns $(entity,relation,entity^{\prime})$. For entities, we convert the mid in KG to the corresponding friendly name. For relations and compound nodes, we keep their original forms in KG. Specifically, we ask LLMs to pick the knowledge triples used to answer $Q$ in a CoT manner. For the example KG instances from “France” in Figure 3, we have 2 knowledge triples “(France, adjoin, compound node), (compound node, country, German)”, based on which, the LLM can reason that the country bordering France” is “German”.
5 Experiments
To comprehensively evaluate the reasoning ability of Readi on large-scale structured environments, we experiment on two complex tasks, KGQA and TableQA (implementation in Appendix A).
5.1 Datasets
We evaluate on three KGQA and two TableQA datasets. Detailed statistics are in Appendix C.
WebQuestionsSP (WebQSP) Yih et al. (2016) contains KGQA questions from Google query logs with up to 2-hop reasoning on Freebase, mostly requiring a single-constrained reasoning path.
ComplexWebQuestions (CWQ) Talmor and Berant (2018) is a complex KGQA benchmark, challenging for up to 4-hop reasoning on Freebase, with 55% multi-constrained questions.
MetaQA Zhang et al. (2018) includes KGQA questions from movie domain, with 3 sub-sets for 3 levels of difficulty, denoted as MQA-1H, MQA-2H and MQA-3H.
WikiTableQuestions (WTQ) Pasupat and Liang (2015) contains questions over 421 tables, challenging for complex aggregation operations, e.g., count, argmax, and sorting.
WikiSQL Zhong et al. (2017) is a large-scale complex dataset based on Wikipedia tables, requiring comparison, aggregation and arithmetic operations.
We model all datasets as an information retrieval task, instead of a semantic parsing one.
Methods WebQSP CWQ MQA-1H MQA-2H MQA-3H
Training-based Method
EmbedKGQA Saxena et al. (2020) 66.6 - 97.5 98.8 94.8
NSM He et al. (2021) 67.7 47.6 97.1 $\underline{99.9}$ 98.9
TransferNet Shi et al. (2021) 71.4 48.6 97.5 $\textbf{100}^{\star}$ $\textbf{100}^{\star}$
SR+NSM+E2E Zhang et al. (2022) 69.5 49.3 - - -
UniKGQA Jiang et al. (2023c) 75.1 50.7 97.5 99.0 99.1
ReasoningLM Jiang et al. (2023b) 78.5 $\textbf{69.0}^{\star}$ 96.5 98.3 92.7
RoG Luo et al. (2024) $\textbf{85.7}^{\star}$ 62.6 - - 84.8
Inference-based Method
Davinci-003 Ouyang et al. (2022) 48.7 - 52.1 25.3 42.5
GPT3.5 OpenAI (2022) 65.7 44.7 61.9 31.0 43.2
GPT4 OpenAI (2023) 70.7 52.1 71.8 52.5 49.2
AgentBench Liu et al. (2024b) 47.8 24.8 - - -
StructGPT Jiang et al. (2023a) 69.6 - 97.1 97.3 87.0
Readi-GPT3.5 74.3 55.6 $\underline{98.4}$ 99.9 99.4
Readi-GPT4 78.7 67.0 $\textbf{98.5}^{\star}$ 99.9 99.2
Table 1: QA performance (Hit@1) of Readi on KGQA datasets. Results of GPT3.5, GPT4 OpenAI (2023) and AgentBench Liu et al. (2024b) are run by ourselves, others are from the origin paper. Bold and underline fonts denotes the best and second-best for two types of methods, respectively. $\star$ denotes the overall state-of-the-art result.
5.2 Baselines
KGQA Baselines. Training-based methods fine-tune pre-tained language models (PLMs): EmbedKGQA Saxena et al. (2020), NSM He et al. (2021), TransferNet Shi et al. (2021), SR+NSM+E2E Zhang et al. (2022), UniKGQA Jiang et al. (2023c), ReasoningLM Jiang et al. (2023b) and RoG Luo et al. (2024). Inference-based methods call LLMs API for inference: Davinci-003 Ouyang et al. (2022), GPT3.5, GPT4 OpenAI (2023), AgentBench Liu et al. (2024b) and StructGPT Jiang et al. (2023a).
For fair comparison, we assume golden topic entities are given. Following Tan et al. (2023), we adopt Hit@1, assessing whether the predicted answer is correct, to evaluate KGQA.
TableQA Baselines. Training-based methods fine-tune PLMs: TAPAS Herzig et al. (2020), UnifiedSKG Xie et al. (2022), TAPEX Liu et al. (2022). Inference-based methods call LLMs API for inference: Davinci-003 Ouyang et al. (2022), GPT3.5, GPT4 OpenAI (2023), StructGPT Jiang et al. (2023a). Note that all Inference-based baselines regard the datasets as an information retrieval task.
We adopt denotation accuracy, assessing whether the predicted answer exactly matches the golden, to evaluate TableQA. Detailed descriptions of each baseline are in Appendix D.
5.3 Implementation Details
We adopt gpt-3.5-turbo OpenAI (2022) and GPT4 OpenAI (2023) as LLMs, denoted as Readi-GPT3.5 and Readi-GPT4. Temperature is 0.3 for all modules. For relation-binding, we deploy a Pyserini as a hybrid searcher with BM25 and Contriever Izacard et al. (2022). For each relation generated by LLMs, we retrieve top 5 similar relations on Freebase Bollacker et al. (2008). For instantiation, we deploy a Virtuoso server following the instructions [^2] https://github.com/dki-lab/Freebase-Setup. Prompt lists are in Appendix F.
5.4 Main Results
Results for KGQA As shown in Table 1, Overall, Readi significantly outperforms all inference-based methods, the vanilla LLMs and most training-based methods on all datasets, demonstrating the effectiveness of Readi. Compared with inference-based methods, Readi substantially boosts the results of the vanilla LLM (by 8.6% on WebQSP and 14.9% on CWQ), demonstrating that Readi enables LLMs to practically interact with structured environments. Moreover, Readi-GPT3.5 already significantly surpasses state-of-the-art results (by 4.7% on WebQSP and 12.4% on MQA-3H), and Readi-GPT4 further enhances the performance. With fewer LLM-calls, our directly-generated reasoning path and editing framework still achieves better performance. Compared with training-based methods, without large-scale supervision and the cost of beam search, Readi achieves comparable performance (e.g. 67.0 Hit@1 on CWQ) just by some demonstration examples, showing the faithfulness of Readi. Additionally, Readi sets new state-of-the-art results MQA-1H, illustrating the effectiveness of the editing module which refines the reasoning path if it goes wrong.
Methods WTQ WikiSQL
Training-based Method
TAPAS 48.8 83.6
UnifiedSKG (T5-3B) 49.3 86.0
TAPEX 57.5 89.5
Inference-based Method
Davinci-003 34.8 49.1
GPT3.5 55.8 59.8
GPT4 57.0 59.9
StructGPT 52.2 65.6
Readi-GPT3.5 61.7 66.2
Readi-GPT4 61.3 66.0
Table 2: QA performance (denotation accuracy) of Readi on TableQA datasets. Bold fonts denotes the best results for Training-based and Inference-based methods.
Variance of Readi Answer Coverage Rate (AC) QA Performance (Hit@1)
Corrupt Empty GPT3.5 GPT4 Corrupt Empty GPT3.5 GPT4
w/o edit - - 56.7 62.7 - - 51.0 57.2
w/ edit by GPT3.5 54.0 56.4 62.5 64.3 57.3 58.5 58.7 58.5
w/ edit by GPT4 55.6 63.9 68.6 65.8 58.2 59.9 58.1 59.3
Table 3: Answer Coverage Rate (AC) and QA Performance (Hit@1) of variance of Readi (GPT3.5 as reasoning module). Each row is by different initial path. Corrupt means corrupted path with some randomly-sampled relations. Empty means empty path. w/o edit means we remove the editing module and only leverage the initial reasoning path.
Results for TableQA We also experiment on TableQA scenario which requires multi-hop reasoning over table to show the generalizability of Readi. As shown in Table 2, overall, Readi outperforms all inference-based baselines and most training-based methods, setting state-of-the-art results on WTQ, demonstrating the effectiveness of our framework. Compared with inference-based methods, Readi surpasses previous state-of-the-art methods by 9.5% and the vanilla LLM by 5.9% on WTQ, demonstrating that LLMs perform well with directly generated reasoning path and some necessary editing. Compared with training-based methods, Readi is significantly superior on WTQ, while trailing behind on WikiSQL, without the cost of massive annotations. This may be due to the i.i.d. distribution between the training and testing sets of WikiSQL, which favors the results of fully-trained methods. Interestingly, Readi-GPT4 lags a little bit behind Readi-GPT3.5 on both datasets, which may be due to the information retrieval manner that asks LLMs to output answer directly based on retrieved table items, relying heavily on mathematical reasoning ability of LLMs, where GPT3.5 already plays on par with GPT4.
6 Analysis
We further analyze Readi’s modules, reasoning path and efficiency on 1000 test samples of CWQ. For fairness, we base all evaluation on our instantiation and reasoning modules.
6.1 Ablation Study
Effectiveness of reasoning path generation and editing modules. As shown in Table 3, we adpot the answer coverage rate (AC, rate of instances containing the answer) and QA performance (Hit@1) for illustration. We also test editing module with a corrupt path and an empty path. First, Readi establishes a plug-and-play nature for both modules, showing their effectiveness. With only an initial path (we instantiate the path and maintain the longest path if it goes wrong), Readi reaches comparable results (1st row). With an empty initial path (“Empty” columns), the editing modules still yield competitive results closer to the full Readi. Second, generally, higher capacity of LLMs leads to better results for both modules. This meets our expectation, suggesting future improvements of Readi. Third, Readi shows robustness for reasoning path editing, performing well even with an empty or even a corrupt path (“Corrupt” and “Empty” columns).
6.2 Reasoning Path Analysis
We compare reasoning path of Readi with representative fine-tuned methods, i.e., Subgraph Retrieval (SR) Zhang et al. (2022) trains an encoder to retrieve relations to build a path, RoG Luo et al. (2024) tunes a Llama 2 Touvron et al. (2023) to generate a path.
Methods Graph Quality QA Perf.
AC #RK Hit@1
SR
   - beam size 1 58.4 26.3 50.9
   - beam size 3 67.2 47.1 54.6
RoG
   - beam size 1 57.0 69.5 52.2
   - beam size 3 77.5 170.1 57.3
Readi initial path
   - GPT3.5 56.7 134.6 51.0
   - GPT4 62.7 101.4 57.2
Readi full
   - GPT3.5 62.5 93.7 58.7
   - GPT4 71.8 121.5 59.3
Table 4: Reasoning path evaluation of Readiand compared methods. AC and #RK denotes answer coverage rate and number of retrieved knowledge, respectively.
How well Readi’s reasoning path is? We adopt answer coverage rate (AC) and number of retrieved knowledge (#RK) as the quality of graph. Ideally, the higher AC and lower #RK, the better. Also, we report Hit@1 as QA performance, shown in Table 4. First, Readi’s initial path is comparable with fine-tuned methods, with GPT4 surpassing them by a large margin (5.0% and 6.3% Hit@1 than RoG and SR, respectively), showing the effectiveness of our reasoning path. Second, with some necessary editing, Readi obtain substantially higher AC (while just a little higher #RK with GPT4 and even lower #RK with GPT3.5) and significantly higher QA performance than fine-tuned methods. Third, with wider beam size, fine-tuned methods gain higher AC, yet a drastically growing knowledge size, and lower Hit@1 than Readi, showing our superiority.
Extensive features of Readi’s reasoning path. The quality of reasoning path is multi-dimensional. To further show insights of reasoning path by LLMs, we design some metrics, including the instantiation progress, error types, etc. Also, we compare with SR Zhang et al. (2022), RoG Luo et al. (2024) and Golden (outermost in Figure 4). More detailed analysis are in Appendix E.
Figure 4: Extensive features of Readi’s reasoning path, compared with fine-tuned methods and Golden.
Insights driven by Figure 4 are three-folded: 1) Readi’s initial path already achieves better in QA results. More than a half (60%) of Readi-GPT4’s initial paths can be fully instantiated (LLMs called only once), showing that Readi effectively unleash the intrinsic planning ability of LLMs. For each path, a large proportion (averagely 74%) can be instantiated, indicating the possibility to safe LLM-calls. 2) With necessary editing, Readi gets significantly closer to golden (from green to red in Figure 4), exceeding compared methods in many metrics, especially the QA results. 3) Interestingly, the path by LLMs establish a human-like nature, for humans tend to get stuck at points they have never seen in real world. For compound node ending rate (special for KG), fine-tuned methods are close to golden, showing that they memorize the structured information. However, they trail behind behind Readi on QA results, demonstrating that unfaithfulness still exits for fine-tuned methods.
Question 1: What is the name of the money used in the country the Peruvian Paso breed originated?
Initial Reasoning Path: (from “Peruvian Paso”)
[…]
Error Message: irrelevant relation “biology…breeds”.
Candidates: biology.breed.originated_in, …
Edited Reasoning Path: (from “Peruvian Paso”)
[…]
Question 2: What to see in the country that has Gozo?
Initial Reasoning Path: (from “Gozo”)
[…]
Error Message: irrelevant relation “location…attractions”.
Contexts: Gozo ${location.location.containedby}$ Melta
Candidates: travel.travel_destination.tourist_attractions, …
Edited Reasoning Path: (from “Gozo”)
[…]
Table 5: Cases of Readi’s reasoning path editing.
6.3 Efficiency Evaluation
How many LLM-calls do we need? Due to unexpectable nature of LLMs output, we cannot give an exact number of LLM-calls for Readi. Instead, we present the distribution of number of LLM-calls for editing in Figure 5 and some exemplar cases in Table 5. Note that there are at least 2-hop and up to 4-hop reasoning required for CWQ (theoretically 4-8 LLM-calls for iterative interaction). The instantiation success rate and average instantiate progress in Figure 4 already demonstrate that Readi can save a bunch of LLM-calls. In Figure 5, most of the time the initial reasoning path is already instantiable, no need for more LLM-calls. Averagely, the LLM is called 1.55 times to edit the path, saving more invocation than iterative interaction. Also, cases in Table 5 show that, with the necessary editing, LLMs can correct previous path based on some error messages during instantiation.
Figure 5: Distribution of number of LLM-Call for reasoning path editing of Readi-GPT4.
7 Conclusion
In this paper, we propose a novel framework Readi where LLMs can reason over structured environments both efficiently and faithfully. In Readi, LLMs initially generate a reasoning path which is then instantiated on environments to facilitate faithful reasoning. Path editing is triggered only if the instantiation gets stuck. We showcase the implementation of Readi on knowledge graph, evaluate the effectiveness on KGQA and TableQA and analyze extensive features of Readi. Our work also shed lights on practically interaction between natural language and structured environments, where LLMs plays a crucial role to bridge the gap.
8 Limitations
Although Readi achieves strong performance in complex reasoning task over structured environments, e.g., knowledge graph and tables, there are still some limitations of our method. First, we leverage just 2 LLMs as backbone to evaluate Readi. Therefore, more experiments can be done to test performance of Readifor other LLMs (including the inference-based and training-based). Second, we adopt reasoning path as a representation of natural language utterance and model the task as a information-retrieval one. An interesting and more precise direction is semantic parsing, e.g., text2SQL, which we will leave for future work. Moreover, we can also test Readi on question answering over database or other reasoning fields. Third, the instantiation in our current implementation of KGQA is intuitive but may induce many queries for SPARQL engine if encountered some “big” entities. For example, when constraining “locations in France”, we may obtain hundreds of instances, for which we just set a threshold to restrict the number of current entities. This can be improved by meticulous query design. Forth, our proposal to fully harness intrinsic planning ability of LLMs performs well and saves some LLM-calls in experiments, which does not mean we should always depend on LLMs plans in every other task. And although the initial path by LLMs can be well instantiated, this just means the path by LLMs corresponds to some instances on KG, but not ensures the path exactly matches with the constraints in the question. Finally, by efficiency, we mean number of LLM-calls, or how LLMs can practically interact with structured environments. A promising direction is to reach for time and cost efficiency.
Appendix A Implementation of Readi for TableQA
Name in English Name in Turkish Area ($km^{2}$) Depth Location (Districts and/or provinces)
Lake Van Van Gölü 3755 km2 3755 km2 Van, Bitlis
Lake Tuz Tuz Gölü 1500 km2 2 m Aksaray, Ankara, Konya
Lake Palas Tuzla Palas Tuzla Gölü 106 km2 15 m Palas/Kayseri
Table 6: An example from WTQ Pasupat and Liang (2015), with the question “which is deeper, lake tuz or lake palas tuzla?”.
Algorithm 1 again demonstrates Readi’s framework. Our implementation of Readi for TableQA strictly sticks to Algorithm 1, and is just simpler than KGQA. The main difference between TableQA and KGQA is the reasoning path (mentioned in Section 3). Reasoning path for TableQA starts from a given table, to some columns and then to some rows, in order to constrain specific items in the table to answer the question. A sample of WTQ Pasupat and Liang (2015) dataset is shown in Table 6. The reasoning path of such question is from this table to column“Name in English” and “Depth”, and then to row items whose “Name in English” is “Lake Tuz” and “Lake Palas Tuzla”.
For reasoning path generation, we ask LLMs to generate a reasoning path for a given question and some table descriptions (including the header and a randomly-sampled example row), by in-context learning. Specifically, we require the LLMs to pick up at least to rows for complex arithmetic and aggregation operations in TableQA datasets.
Data: task $Q$, entity set $E$, environment $\mathcal{G}$, max edit time $T$
1 $t\leftarrow 0$
2 $P_{0}\leftarrow\textbf{Reasoning\_Path\_Generation}(Q,E)$
3 while $t\leq T$ do
4       $I_{\mathcal{G}},Err_{msg}\leftarrow\textbf{Instantiate}(P_{t},\mathcal{G})$
5       if $Err_{msg}\neq\emptyset$ then
6             $P_{t+1}\leftarrow\textbf{Edit\_Path}(Err_{msg},P_{t},Q)$
7             $t\leftarrow t+1$
8      else
9             Go to Line 11
10       end if
11
12 end while
$Instances_{\mathcal{G}}\leftarrow\textbf{Merge\_Results}(I_{\mathcal{G}})$
Algorithm 1 Readi
For reasoning path instantiation, since schemas of a table is not as massive as a KG, we don’t need a relation binding module to retrieve candidate relations and connect them according to the reasoning path. For a table, instead, we first filter out columns and then filter out the rows in the reasoning path. If the selected columns or rows are not in the table, this is the "necessary" time to call LLMs to edit the path.
For reasoning path editing, similar to KGQA, we show the currently instantiated columns, and then give out candidate columns where the instantiation goes wrong, as error messages. Note that most questions of WTQ Pasupat and Liang (2015) and WikiSQL Zhong et al. (2017), so we just consider wrong columns as error. If a selected row in the reasoning path cannot be instantiated, we return all rows in the columns. Then we ask the LLMs to edit its reasoning path.
The QA reasoning is quite similar to KGQA, we concatenate the question and retrieved row items and ask the LLMs to answer the question.
Appendix B Detailed Implementation of Readi for KGQA
The example prompts of reasoning path generation and reasoning path editing are in Appendix F. We demonstrate some details here.
B.1 Reasoning Path Instantiation Details
In this part, we demonstrate the detailed path-connecting part mentioned in Section 4.3. Assume that we have a reasoning path starting from an entity, with $m$ relations $R$. All relations are binded to candidates $\hat{R}$. For the example in Figure 3, we have a reasoning path “[Nijmegen] serve_airport $\rightarrow$ contain” and have “$r_{1}:$ serve_airport” binded to “$\hat{r}_{1}:$ airport, terminal, serving port” in KG, and similarly have “$r_{2}:\leavevmode\nobreak$contain” binded to “$\hat{r}_{2}:\leavevmode\nobreak$contained by, contains, place of”. The purpose if this part is to check if there exists an instance in KG where the starting point is entity “Nijmegen”, then connect to some relations in $\hat{r}_{1}$, and then connect to some relations in $\hat{r}_{2}$.
The path-connecting is essentially an Breadth-first search (BFS) algorithm, with a queue $que$ containing possible entities connected to current relations. Firstly, we push the starting point (e.g., “Nijmegen”) into $que$. Second, at each time, we check if current candidate relations (e.g.,adjoin, terminal, serving port”) can connect to current head of $que$. If so, we successfully instantiate “serve_airport” in KG, and we can obtain some entities which are “serve_airport” of “Nijmegen” (e.g., “WZ air.” and “NTA.”). Then, we add these entities into the queue, and then match them with the rest relations.
[…]
B.2 Reasoning Path Editing Details
Section 4.4 has covered the basic process of reasoning path editing. Here are some implementation details for preparation step, the purpose is to help LLMs identify the error position of previous reasoning path and edit it according to some tips. For the half-way done instances in the Error Messages, we use “compound node” to indicate all cvt nodes in Freebase  Bollacker et al. (2008), and we just sample some instances of each instantiated relation to showcase the path instances in KG of the previous path. The loss of information from sampling is minor because the messages just inform the LLMs that the relations in the reasoning path can be grounded to some certain instances on KG. For candidates, we use a threshold to constrain the size, if there are too many of them, we filter out using similarity search (same as relation-binding module) according to the original question.
[…]
Appendix C Datasets
Statistics of datasets are shown in Table C. We experiment Readi on test sets of all datasets. We adopt Hit@1 and denotation accuracy for KGQA and TableQA, respectively. We evaluate Readi on 3 KGQA and 2 TableQA datasets.
WebQuestionsSP (WebQSP) Yih et al. (2016) contains KGQA questions from Google query logs with up to 2-hop reasoning path on Freebase, mostly requiring a list structured reasoning path.
ComplexWebQuestions (CWQ) Talmor and Berant (2018) is a complex KGQA benchmark, challenging for up to 4-hop reasoning path on Freebase, with 55% multi-constrained questions (a tree structured reasoning path).
MetaQA Zhang et al. (2018) includes KGQA questions from movie domain , with 3 sub-sets for 3 levels of difficulty, denoted as MQA-1H, MQA-2H and MQA-3H.
WikiTableQuestions (WTQ) Pasupat and Liang (2015) contains questions over 421 tables challenging for complex aggregation operations, e.g., count, argmax, and sorting.
WikiSQL Zhong et al. (2017) contains complex question based on Wikipedia tables, requiring comparison, aggregation and arithmetic operations.
Dataset Training Dev Test
KGQA Dataset
WebQSP 3,098 - 1,639
CWQ 27,639 3,519 3,531
MQA-1H 96,106 9,992 9,947
MQA-2H 118,980 14,872 14,872
MQA-3H 114,196 14,274 14,274
TableQA Dataset
WTQ 11,321 2,831 4,344
WikiSQL 56,355 8,421 15,878
Table 7: Statistics of experiment datasets.
Appendix D Baselines
KGQA baselines. Training-based methods:
  • EmbedKGQA Saxena et al. (2020) adopts an encoder to retrieve relevant entities and generate the answer.
  • NSM He et al. (2021) adapts neural state machine for KGQA and rank entities in a retrieved subgraph.
  • TransferNet Shi et al. (2021) adopts a transparent framework to rank entities according to different parts of a question in a subgraph.
  • SR+NSM+E2E Zhang et al. (2022) trains an encoder to retrieve relevant relations and build a path from retrieved relations.
  • UniKGQA Jiang et al. (2023c) retrieves a sub-graph and rank the schemas in a unified way.
  • ReasoningLM Jiang et al. (2023b) designs an entity encoding and training framework to rank entities in a sub-graph.
  • RoG Luo et al. (2024) trains a LLama 2 Touvron et al. (2023) to firstly generate a path, second ground this path to the knowledge graph, and then generate answer based on the grounded graph.
Inference-based methods:
  • Davinci-003 Ouyang et al. (2022), GPT3.5 and GPT4 OpenAI (2023) are based on LLM-APIs. We adopt few-shot in-context learning to ask the model to output the answer of question in order to test the models inherent knowledge of datasets.
  • AgentBench Liu et al. (2024b) is an agent-based method asking LLMs to call tools based on tool-discription, history and observations.
  • StructGPT Jiang et al. (2023a) requires LLMs to iterative pick up relations and entities based on current returned candidates.
TableQA baselines. Training-based methods:
  • TAPAS Herzig et al. (2020) predicts denotation by selecting table cells and optionally applys an aggregation operator to the selection.
  • UnifiedSKG Xie et al. (2022) sequentializes the table and tunes a T5-3B model to answer the question.
  • TAPEX Liu et al. (2022) guides language models to mimic a SQL executor.
Inference-based methods:
  • Davinci-003 Ouyang et al. (2022), GPT3.5, GPT4 OpenAI (2023) are based on LLM-APIs. We adopt few-shot in-context learning to ask the model to output the answer, based on the question and the entire table.
  • StructGPT Jiang et al. (2023a) iteratively filter out columns and rows of table and ask the LLMs to answer the question.
Note that all Inference-based baselines are information retrieval ones.
  Metrics LPP LIP AIP ISR CER
  Golden Path 3.2 3.2 1.0 1.0 0
  SR Zhang et al. (2022) 3.7 3.3 0.88 0.86 0.01
  RoG Luo et al. (2024) 2.6 1.4 0.55 0.50 0.02
  GPT3.5 4.3 2.5 0.64 0.46 0.49
      + editing 3.3 2.8 0.86 0.80 0.22
  ChatGPT4 3.4 2.5 0.74 0.60 0.45
      + editing 3.6 3.2 0.89 0.84 0.25
Table 8: Extensive features of Readi-GPT3.5 and Readi-ChatGPT4’s reasoning path, compared with Golden reasoning path. LPP, LGP, AGP, GSR, IER means length of predict path, length of instantiated path, average instantiation progress, grounding success rate and ending with intermediate nodes rate, respectively.
Figure 6: Extensive analysis of Readi-GPT3.5 reasoning path, compared with Golden reasoning path.
Figure 7: Features of Readi’s reasoning path, compared with fine-tuned methods and Golden.
Figure 8: Distribution of number of LLM-Call for reasoning path editing of Readi-GPT3.5
Appendix E Detailed Features of Reasoning Path
Note that quality of reasoning path is multi-dimensional. To further show some insights of reasoning path by LLMs, we design some metrics, including the instantiation progress, error types, etc. We introduce average Length of predict path (LPP), average length of instantiated path (LIP), average instantiation progress (AIP, average of LIP/LPP for each question), instantiation success rate (ISR) and Compound nodes Ending Rate (CNR), where all length mentioned above refers to number of relations on path. Also, we compare with path by SR Zhang et al. (2022) and RoG Luo et al. (2024) (beam = 1). The Golden Path is obtained by extracting relations from golden logical forms in the dataset, just used for analysis.
The absolute performance of different metrics of Readi are shown in Table 8. Also, Figure 6 demonstrates some instantiate progress of Readi-GPT3.5 and Figure 7 shows some relative results of different metrics, compared with fine-tuned methods and the golden path. Additionally, we show the distribution of number of LLM-calls for reasoning path editing in Figure 8.
It is shown that the initial reasoning path by Readi is already comparable with fine-tuned ones. With some necessary editing, Readi significantly gets closer to golden. In addition, not awared of the environments, a large number of Readi’s initial path (49%) get stuck at compound nodes, while fine-tuned methods well memorize the schemas in structured environments. However, with error messages, Readi substantially alleviates this problems and reaches higher QA performance than fine-tuned methods, again demonstrating that unfaithfulness still exits even with large scale fine-tuning.
Prompts for reasoning path generation
Given a question and some Topic Entities in the Question, output possible freebase Relation Paths starting from each Topic Entities in order to answer the question.
Demonstration Example
Question: Find the person who said “Taste cannot be controlled by law”, where did this person die from?
Topic Entities: [“Taste cannot be controlled by law”]
Thought: There is only one topic entity, the answer is constrained by one path. For, the path from “Taste cannot be controlled by law”, firstly, it should cover the person quote it. Second, it should cover the place where the person died.
[…]
Prompts for reasoning path editing
Task: Given an Inital Path and some feedback information of a Question, please correct the initial path.
Demonstration Example
Question: The movie featured Miley Cyrus and was produced by Tobin Armbrust?
[…]
1. <compound node> in the end.
2. relation "film.film.producer" not instantiated.
Instantiation Context
Instantiate Paths: Miley Cyrus $\rightarrow$ film.actor.film $\rightarrow$ <compound node>
Candidate Relations
[[[[’film.director.film’, ’film.performance.film’, …]]]]
Corrected Path
Goal: The Initial Path starts from Miley Cyrus, which should cover the movies featured by Miley Cyrus.
Thought: In Instantiate Paths I know that Miley Cyrus acts some films, described by a compound node. In candidates, I find "film.performance.film" most relevant to get the films. Meanwhile, "film.film.producer" is not relevant to my Goal.
[…]
Prompts for QA reasoning
Given a question and the associated retrieved knowledge graph triplets (entity, relation, entity), you are asked to answer the question with these triplets. If the given knowledge triples is not enough or missing, you can use your own knowledge. Use {} to enclose the answer! Please think step by step.
Demonstration Example
Q: The artist nominated for The Long Winter lived where?
Knowledge Triplets:
(The Long Winter, book.written_work.author, Laura Ingalls Wilder), (Laura Ingalls Wilder, people.person.places_lived, m.28e5697), (m.28e5697, people.place_lived.location, De Smet)
A: First, based on (The Long Winter, book.written_work.author, Laura Ingalls Wilder), the author of The Long Winter is Laura Ingalls Wilder. Second, based on (Laura Ingalls Wilder, people.person.places_lived, m.28e5697), (m.28e5697, people.place_lived.location, De Smet), Laura Ingalls Wilder lived in De Smet. So, the answer is {De Smet}.
Table 9: Detailed prompts for modules of Readi.
Appendix F Prompt List
Detailed prompts for each module of Readi is in Table 9. For demonstrations, on CWQ Talmor and Berant (2018) and WebQSP Yih et al. (2016) we adopt 6 shots for reasoning path generation and 5 shots for other modules. For MQAZhang et al. (2018), the shot number for reasoning path generation follows Li et al. (2023) and we adopt only 3 shots for editing and reasoning (we donnot design a reasoning module for MQA). For WTQ Pasupat and Liang (2015) and WikiSQL Zhong et al. (2017), we use 7, 2, 7 shots for generation, editing and reasoning, respectively.